A 0.170 mole quantity of NiCl2 is added to a liter of 1.20 M NH3 solution. What is the concentration of Ni2+ ions at equilibrium? Assume the formation constant of Ni(NH3)2+6 is 5.5×108 .
1
Expert's answer
2020-04-23T13:09:02-0400

Solution:

1) 0.170 mole of NiCl2 is added to a liter of 1.20 M NH3 solution.

Therefore, the concentration of NiCl2 is 0.170 mol / 1L = 0.170 M.

Co(Ni2+) = C(NiCl2) = 0.170 M.

2) Concentration of the NH3 solution is 1.20 M.

The number of moles of ammonia is 1.20 mol.


The balanced chemical equation is:

Ni2++6NH3 ⇒ [Ni(NH3)6]2+

The formation constant of the reaction is Kf = 5.5×108.

The formation constant of the reaction is very high so we can say that almost the reactants are consumed up to the maximum level.

According to the chemical equation: 1 mole of Ni2+ reacts with 6 moles of NH3.

Thus, we have 0.170 mol of Ni2+ reacts with 6 × 0.170 = 1.02 moles of NH3.

Amount of NH3 remain in the solution is n = (1.20 − 1.02) = 0.18 mol.

Let's write the ICE (Initial, Change, Equilibrioum) table:




5.5×108 * x * 0.186 = 0.17;

x = [Ni2+] = 9.09×10-6 M.

[Ni2+] = 9.1×10-6 M.


Answer: [Ni2+] = 9.1×10-6 M.